OpenStudy (sparklestaraa):

help need clarification on the topic of Calculating Variance. I have one step of the process I'm stuck on screenshots below

OpenStudy (sparklestaraa):

working on part 1 and the #'s i wrote are for Class A. Im wondering step 2 its confusing i did the whole subtracting mean but the squared difference part is confusing is square difference just the square root and the # for ex. 9 the squared difference is 81? or 3?

OpenStudy (sparklestaraa):

if this helps any heres the question

OpenStudy (sparklestaraa):

@InExileWeTrust

OpenStudy (cruffo):

Let \(x_1, x_2, ..., x_n\) be the n data values. Mean: \( \bar{x} = \dfrac{x_1 + x_2 + ... + x_n}{n}\) Variance = \(\dfrac{(x_1 - \bar{x})^2 + (x_2 - \bar{x})^2 + ... + (x_n - \bar{x})^2}{n}\)

OpenStudy (anonymous):

If the mean is \(\mu\) then the squared difference is \((x-\mu)^2\).

OpenStudy (perl):

I got a different mean for class A ,

OpenStudy (sparklestaraa):

ok if i understand right the numbers i have there are right then right? now all i have to do is ^2 each?

OpenStudy (amistre64):

i just downed openoffice today ... excited to use its excel clone

OpenStudy (cruffo):

Hmmm... I got a Class A average of 79 as well. So, this is how variance would look worked out: (80 - 79)^2 = (1)^2 = 1 (85 - 79)^2 = (6)^2 = 36 (87 - 79)^2 = (8)^2 = 64 ... (62 - 79)^2 = (-17)^2 = 289 Then add up all those results 1 + 36 + 64 + ... +289 = 4558 and divide by 15 4558/15 = 303.866666.......

OpenStudy (amistre64):

it works like a remember it did :)

OpenStudy (sparklestaraa):

Thanks :D i think i understand I'm going to ttry it again and i will come back here if i don't get it